# Investigating Adversarial Attacks in Software Analytics via Machine Learning Explainability

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; ; price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Machine learning explainabilityccs: Just-in-Time (JIT) defect prediction, Code clone detection.
1. Introduction
@@
In recent years, machine learning (ML) models have been extensively applied in software analytics tasks such as code clone detection (Nafi et al., 2019; Feng et al., 2024; Hu et al., 2022; Wu et al., 2022), API recommendation (Moreno et al., 2015; Nguyen et al., 2019), automatic code summarization (McBurney and McMillan, 2015; Zhu and Pan, 2019), malware detection (Sahs and Khan, 2012; Gavriluţ et al., 2009), code completion (Das and Shah, 2015; Svyatkovskiy et al., 2021), Just-in-Time (JIT) defect prediction (Catolino et al., 2019; Yatish et al., 2019; Kamei et al., 2012), and source code authorship attribution (Alsulami et al., 2017; Ullah et al., 2019). However, research shows that despite the effectiveness of the state-of-the-art ML models, they are vulnerable to adversarial attacks (Szegedy et al., 2013; Liu et al., 2021; Quiring et al., 2019; Chen et al., 2017a; Bielik and Vechev, 2020; Springer et al., 2020; Jha and Reddy, 2023; Yang et al., 2022; Zhou et al., 2022; Zhang et al., 2020; Severi et al., 2021). For example, carefully crafted data from large open-source software repositories in an API recommender system can lead to malicious API calls by developers (Nguyen et al., 2021). As a result, researchers have proposed various techniques for generating adversarial examples to evaluate the robustness of the ML models in software analytics tasks (Zeng et al., 2022; Du et al., 2023; Nguyen et al., 2021; T. Nguyen et al., 2021; Suciu et al., 2019; Liu et al., 2019, 2021; Quiring et al., 2019; Bielik and Vechev, 2020; Springer et al., 2020; Jha and Reddy, 2023; Yang et al., 2022; Zhou et al., 2022; Zhang et al., 2020).
@@
Adversarial attacks generate perturbations to deceive and manipulate ML model predictions, potentially leading to system failures in real-life applications. For example, failing to predict JIT defect introducing commit using ML models under adversarial attacks can significantly impact software development and maintenance processes and optimize the allocation of limited software quality assurance (SQA) resources (Pornprasit et al., 2021; Jiarpakdee et al., 2020). It can also affect the quality maintenance and execution of exhaustive code review activities for all incoming commits with limited SQA resources in modern code review (McIntosh et al., 2014; Bosu et al., 2015). Consequently, the overall software development and maintenance process may experience significant delays, resulting in financial losses and wasted time, as the underlying ML models struggle to make accurate decisions in the face of adversarial attacks (Kamei et al., 2012; Catolino et al., 2019). Furthermore, attacks on ML models used in code review comment classification tasks can affect automatic support for improving non-useful comments in code review (Rahman et al., 2017). Mondal et al. (Mondal et al., 2019) empirically showed that code clones are directly related to bugs, and the intensity of bug propagation through code cloning is significant. In addition, as software systems evolve and the demand for code grows, the occurrence of code clones may pose a significant threat to vulnerability propagation within these systems (Feng et al., 2024). Thus, failure to detect code clones accurately under adversarial attacks can lead to the development of buggy software, resulting in technical debt and a huge loss (\$$\$3.61$ per line of code) in monetary value (Guo et al., 2016; McGraw-Hill Book Co, 2012). Therefore, having a thorough understanding of the performance and resilience of ML models in various software analytics tasks regarding input manipulation is crucial for developing robust ML models.
@@
Most research on adversarial attacks in software analytics tasks focuses on deep learning (DL) models trained on the source code (Zeng et al., 2022; Du et al., 2023; Nguyen et al., 2021; T. Nguyen et al., 2021; Liu et al., 2019, 2021; Quiring et al., 2019; Bielik and Vechev, 2020; Springer et al., 2020; Jha and Reddy, 2023; Yang et al., 2022; Zhou et al., 2022; Zhang et al., 2020). Furthermore, the attack approach targets the input-space (e.g., source code) when generating adversarial examples. However, in other software analytics tasks such as JIT defect prediction, clone detection, and code review comment classification, many classical  [^1] In this study, unless otherwise stated, ML models refer to classical machine learning models, not deep learning models. ML models (not DL) have been extensively used. For example, Feng et al. (Feng et al., 2024), Hu et al. (Hu et al., 2022), and Wu et al. (Wu et al., 2022) demonstrated that ML models (e.g., Random Forest) outperformed several state-of-the-art DL-based approaches (e.g., ASTNN (Zhang et al., 2019), SCDetector (Wu et al., 2020), DeeSIM (Zhao and Huang, 2018), FCCA (Hua et al., 2020), CDLH (Wei and Li, 2017), and TBCNN (Mou et al., 2014)) in software clone detection on the BigCloneBench (Svajlenko et al., 2014) dataset. Moreover, in these tasks, features from the input data are extracted using a pre-trained machine learning model or some other metrics first (Kamei et al., 2012; Catolino et al., 2019; Nafi et al., 2019) and then represented in tabular format. These tabular data are later used to tackle the tasks of interest. Ravid et al. (Shwartz-Ziv and Armon, 2022) demonstrated that ML models trained on tabular data outperformed DL models trained on the same data in various tasks. Therefore, it is essential to assess the robustness of ML models trained on tabular data through feature-space manipulation in software analytics tasks. Adversarial attacks in feature-space have been studied in computer vision (Xu et al., 2020; Inkawhich et al., 2019; Simonetto et al., 2021). However, to the best of our knowledge, this area remains largely unexplored in software analytics tasks. Therefore, our study focuses on bypassing the input-space and instead examines attacks on the feature-space for ML models in software analytics tasks.
@@
Despite significant efforts to advance adversarial attacks targeting DL models trained on source code (Zeng et al., 2022; Du et al., 2023), to the best of our knowledge, adversarial attacks on classical ML models trained on tabular data for various software analytics tasks have not been explored. Moreover, due to fundamental differences in the characteristics of the training datasets and the model’s architecture, the methods proposed for adversarial attacks on DL models cannot be directly applied to ML models trained on tabular data. Therefore, there is an urgent need to propose novel techniques to perform adversarial attacks on ML models trained specifically on tabular data in software analytics tasks. Furthermore, investigating adversarial attacks using machine learning explainability in software analytics tasks remains largely unexplored. Recently, Severi et al. (Severi et al., 2021) proposed explanation-guided backdoor poisoning attacks to assess the robustness of malware classifiers. The primary distinction between Severi et al. (Severi et al., 2021) and our work is that they focused on influencing the ML training process, while our approach is specifically designed to attack the ML inference process. Similarly, Amich et al. (Amich and Eshete, 2021, 2022) employ ML explainability to boost and diagnose evasion attacks on ML models, particularly those trained on image classifier datasets. However, our approach diverges from previous studies (Severi et al., 2021; Amich and Eshete, 2021, 2022) in terms of how we select important features, modify them, and attack feature-space while generating adversarial examples, making our work novel.
@@
This study aims to assess the robustness of ML models in software analytics tasks by generating adversarial examples using ML explainability techniques such as SHapley Additive exPlanation (SHAP) (Lundberg and Lee, 2017), Local Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016), and PyExplainer (Pornprasit et al., 2021). The key contributions of our study are listed below:
@@
  • We comprehensively evaluate the proposed adversarial example generation technique using six datasets and seven ML models. Additionally, we compare our approach with four state-of-the-art adversarial attack techniques (e.g., Zoo (Chen et al., 2017b), Boundary attack (Brendel et al., 2017), PermuteAttack (Hashemi and Fathi, 2020), and HopSkipJump (Chen et al., 2020)) designed for ML models in a model-agnostic way. We also compare our approach with two makeshift tools. Our proposed approach demonstrates promising results compared to the baselines and makeshift tools, showcasing its effectiveness in generating adversarial examples to evaluate the robustness of ML models.
@@
2. Background
This section briefly discusses some basic terminologies related to this study.
Adversarial Attack: When a subtle modification to an original input results in a transformed version resembling the original, we refer to the changed version as an adversarial example. If the modified sample alters prediction made by ML model, it is categorized as an adversarial attack (Szegedy et al., 2013).
Adversarial Robustness: If a ML model can perform well in adversarial attacks and maintain accuracy on adversarial examples, we consider it robust against such attacks.
@@ 2. Background
Machine Learning Explainability: ML explainability refers to the ability to understand and interpret the decisions made by a ML model (Ribeiro et al., 2016). In other words, it allows us to explain why a model makes a particular prediction or decision.
3. Research Methodology
@@
Figure 1. Our proposed workflow for evaluating the robustness of ML models under adversarial attacks. Step 1 includes the dataset selection and preprocessing methods. Step 2 involves training ML models, while Step 3 involves selecting ML explainability techniques. Step 4 then describes how we perform adversarial attacks using ML explainability, and Step 5 investigates the robustness of ML models under adversarial attacks.
3.1. Dataset Selection and Preprocessing
@@
This study aims to evaluate the robustness of ML models by generating adversarial examples through feature-space manipulation. Therefore, dataset selection and preprocessing are vital for enhancing the validity and generalizability of the ML results, ensuring that the models perform optimally before undergoing adversarial attacks. We meticulously select and process the datasets as follows:
3.1.1. Dataset Selection
@@
This study investigates the robustness of ML models against adversarial attacks across various software analytics tasks, including defect prediction, clone detection, and code review comment classification. Additionally, our study shifts the focus of attack from the input-space to the feature-space when generating adversarial examples. Thus, we aim to select software analytics tasks where features extracted from the input (e.g., source code) are utilized to perform the tasks of interest. For example, in cross-language clone detection, various features such as the ‘Number of Variables Declared,’ ‘Total Number of Operators,’ ‘Number of Arguments,’ ‘Number of Expressions,’ ‘Total Number of Operands,’ ‘Number of Loops (for, while),’ ‘Number of Exceptions Thrown,’ ‘Number of Exceptions Referenced,’ and ‘McCabe’s Cyclomatic Complexity’ were extracted first, and ML models were then trained based on these extracted features (Nafi et al., 2019). Consequently, we select six different datasets: three for defect prediction, two for clone detection, and one for code review comment classification, as the targets for adversarial attacks, where features are extracted from the input (e.g., source code) and stored in tabular format.
@@
For JIT defect prediction, we utilize a dataset for cross-project mobile apps (Catolino et al., 2019) and ten defect prediction datasets from various Java projects (Yatish et al., 2019). Additionally, we select the Postgres dataset (Kamei et al., 2012) from six subject systems for the Desktop application. The code review comment dataset (Rahman et al., 2017) comprises 15 features and approximately 1,100 data points for detecting useful comments. Finally, we extend our experimental analysis by incorporating the cross-language clone dataset used by Nafi et al. (Nafi et al., 2019) and the BigCloneBench (BCB) dataset (Svajlenko et al., 2014). The selected cross-project mobile apps dataset contains 14 features and approximately 30,000 data points, while different Java projects have datasets with 65 features and approximately 25,000 data points. The CLCDSA dataset consists of 18 features and approximately 30,000 data points. Feng et al. (Feng et al., 2024) extracted six features from the BCB dataset based on some similarity metrics. In our study, we also use these six features to train the ML models. The preprocessed dataset ended up with six features and approximately 546K samples. Please refer to these articles (Svajlenko et al., 2014; Kamei et al., 2012; Catolino et al., 2019; Nafi et al., 2019; Rahman et al., 2017) for details about the features of each dataset.
3.1.2. Feature Engineering

### v2

@@ Abstract
1 Introduction
@@
In recent years, machine learning (ML) models have been extensively applied in software analytics tasks such as code clone detection [1, 2, 3, 4], API recommendation [5, 6], automatic code summarization [7, 8], malware detection [9, 10], code completion [11, 12], Just-in-Time (JIT) defect prediction [13, 14, 15], and source code authorship attribution [16, 17]. However, research shows that despite the effectiveness of the state-of-the-art ML models, they are vulnerable to adversarial attacks [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. For example, carefully crafted data from large open-source software repositories in an API recommender system can lead to malicious API calls by developers [29]. As a result, researchers have proposed various techniques for generating adversarial examples to evaluate the robustness of the ML models in software analytics tasks [30, 31, 29, 32, 33, 34, 19, 20, 22, 23, 24, 25, 26, 27].
@@
Adversarial attacks generate perturbations to deceive and manipulate ML model predictions, potentially leading to system failures in real-life applications. For example, failing to predict JIT defect introducing commit using ML models under adversarial attacks can significantly impact software development and maintenance processes and optimize the allocation of limited software quality assurance (SQA) resources [35, 36]. It can also affect the quality maintenance and execution of exhaustive code review activities for all incoming commits with limited SQA resources in modern code review [37, 38]. Consequently, the overall software development and maintenance process may experience significant delays, resulting in financial losses and wasted time, as the underlying ML models struggle to make accurate decisions in the face of adversarial attacks [15, 13]. Furthermore, attacks on ML models used in code review comment classification tasks can affect automatic support for improving non-useful comments in code review [39]. Mondal et al. [40] empirically showed that code clones are directly related to bugs, and the intensity of bug propagation through code cloning is significant. In addition, as software systems evolve and the demand for code grows, the occurrence of code clones may pose a significant threat to vulnerability propagation within these systems [2]. Thus, failure to detect code clones accurately under adversarial attacks can lead to the development of buggy software, resulting in technical debt and a huge loss (\$$\$3.61$ per line of code) in monetary value [41, 42]. Therefore, having a thorough understanding of the performance and resilience of ML models in various software analytics tasks regarding input manipulation is crucial for developing robust ML models.
@@
Most research on adversarial attacks in software analytics tasks focuses on deep learning (DL) models trained on the source code [30, 31, 29, 32, 34, 19, 20, 22, 23, 24, 25, 26, 27, 43]. Furthermore, the attack approach targets the input-space (e.g., source code) when generating adversarial examples. However, in other software analytics tasks such as JIT defect prediction, clone detection, and code review comment classification, many classical  [^1] In this study, unless otherwise stated, ML models refer to classical machine learning models, not deep learning models. ML models (not DL) have been extensively used. For example, Feng et al. [2], Hu et al. [3], and Wu et al. [4] demonstrated that ML models (e.g., Random Forest) outperformed several state-of-the-art DL-based approaches (e.g., ASTNN [44], SCDetector [45], DeepSIM [46], FCCA [47], CDLH [48], and TBCNN [49]) in software clone detection on the BigCloneBench [50] dataset. Moreover, in these tasks, features from the input data are extracted using a pre-trained machine learning model or some other metrics first [15, 13, 1] and then represented in tabular format. These tabular data are later used to tackle the tasks of interest. Ravid et al. [51] demonstrated that ML models trained on tabular data outperformed DL models trained on the same data in various tasks. Therefore, it is essential to assess the robustness of ML models trained on tabular data through feature-space manipulation in software analytics tasks. Adversarial attacks in feature-space have been studied in computer vision [52, 53, 54]. However, to the best of our knowledge, this area remains largely unexplored in software analytics tasks. Therefore, our study focuses on bypassing the input-space and instead examines attacks on the feature-space for ML models in software analytics tasks.
@@
Despite significant efforts to advance adversarial attacks targeting DL models trained on source code [30, 31], to the best of our knowledge, adversarial attacks on classical ML models trained on tabular data for various software analytics tasks have not been explored. Moreover, due to fundamental differences in the characteristics of the training datasets and the model’s architecture, the methods proposed for adversarial attacks on DL models cannot be directly applied to ML models trained on tabular data. Therefore, there is an urgent need to propose novel techniques to perform adversarial attacks on ML models trained specifically on tabular data in software analytics tasks. Furthermore, investigating adversarial attacks using machine learning explainability in software analytics tasks remains largely unexplored. Recently, Severi et al. [28] proposed explanation-guided backdoor poisoning attacks to assess the robustness of malware classifiers. The primary distinction between Severi et al. [28] and our work is that they focused on influencing the ML training process, while our approach is specifically designed to attack the ML inference process. Similarly, Amich et al. [55, 56] employ ML explainability to boost and diagnose evasion attacks on ML models, particularly those trained on image classifier datasets. However, our approach diverges from previous studies [28, 55, 56] in terms of how we select important features, modify them, and attack feature-space while generating adversarial examples, making our work novel.
@@
This study aims to assess the robustness of ML models in software analytics tasks by generating adversarial examples using ML explainability techniques such as SHapley Additive exPlanation (SHAP) [57], Local Interpretable Model-agnostic Explanations (LIME) [58], and PyExplainer [35]. The key contributions of our study are listed below:
@@
  • We comprehensively evaluate the proposed adversarial example generation technique using six datasets and seven ML models. Additionally, we compare our approach with four state-of-the-art adversarial attack techniques (e.g., Zoo [59], Boundary attack [60], PermuteAttack [61], and HopSkipJump [62]) designed for ML models in a model-agnostic way. We also compare our approach with two makeshift tools. Our proposed approach demonstrates promising results compared to the baselines and makeshift tools, showcasing its effectiveness in generating adversarial examples to evaluate the robustness of ML models.
@@
2 Background
Adversarial Attacks: Szegedy et al. [18] were the first to introduce the concept of adversarial attacks within the domain of computer vision. Their work demonstrated that minimal, pixel-level perturbations—imperceptible to the human eye—could successfully deceive even state-of-the-art image classifiers. Extending this idea to discrete domains, such as clone detection, a classifier $f:X\rightarrow Y$ aims to determine whether two code snippets are similar. In this context, an adversarial attack may apply small, semantic-preserving transformations—such as variable renaming, dead code insertion, or statement reordering—to produce an adversarial example $x^{\prime}$ from an original input $x$, such that $f(x^{\prime})\neq f(x)$, misleading the model without altering the code’s functionality. Similarly, in just-in-time defect prediction, models classify code changes as either buggy or clean. Adversarial examples in this setting can be created by subtly modifying the code—e.g., inserting dead code or removing redundant lines—while preserving the model’s functionality, thereby causing it to change its prediction.
An adversarial example is defined as a slightly modified version of an original input that remains visually indistinguishable from the original to a human observer. Despite these minimal changes, the modified input is crafted with the specific intent to deceive a model into making an incorrect prediction. Formally, given an input $x$ and a small perturbation $\delta$, the adversarial example $x^{\prime}=x+\delta$ is constructed such that $f(x^{\prime})\neq f(x)$, where $f$ denotes the ML model’s prediction function. When such a perturbation successfully alters the model’s output without altering the input’s actual semantics or functionality, the phenomenon is referred to as an adversarial attack [18].
In practice, the implications of such attacks become clearer when we examine how they can subvert real-world software analytics pipelines. To illustrate this, consider a real-world scenario involving a large software company that employs a machine learning–based just-in-time (JIT) defect prediction model integrated into its continuous integration (CI) pipeline. This model automatically analyzes every submitted code commit and predicts whether it is likely to introduce a bug. Commits predicted as buggy are flagged and blocked for further review, while clean commits are automatically merged. By doing so, the model influences key software engineering practices. This enables software developers to better allocate their limited software quality assurance (SQA) resources by reviewing the highest-risk commits first [13, 15]. It also supports the development of proactive measures aimed at enhancing software quality and preventing the recurrence of issues that may lead to bugs in future releases [37].
Yet, the same automation that improves efficiency may open the door to adversarial vulnerabilities. For example, consider an adversary—such as a malicious developer or a third party—who is aware of this automated screening process and intercepts the input to the machine learning model before committing code changes. The adversary subtly modifies the input in a way that preserves the functionality of the buggy code but disrupts the feature patterns learned by the model. As a result, the model misclassifies the buggy commit as clean, allowing it to bypass the code review process. Consequently, the concurrency bug is deployed to production and remains undetected until it causes a critical system failure under high load. Over time, such adversarial evasions accumulate, contributing to technical debt, complicating future maintenance, increasing debugging costs, and degrading overall software quality [40, 41, 42]. Such attacks are not merely technical exploits—they can serve for deliberate strategic purposes. By causing critical bugs to slip into production, an adversary can significantly damage the reputation of the software company, erode user trust, or even trigger substantial financial losses due to system outages, regulatory penalties, or gain an unfair market advantage. This scenario illustrates how adversarial attacks can have far-reaching consequences that extend beyond model accuracy, affecting software reliability, business operations, and stakeholder trust. Thus, it underscores the importance of assessing model robustness against such attacks prior to deployment in real-world software engineering tasks.
Adversarial Robustness: Adversarial robustness refers to a model’s ability to maintain its predictive performance when subjected to adversarial perturbations—small, intentionally crafted modifications to input data designed to mislead the model into making incorrect predictions. A model is considered adversarially robust if its output remains stable and consistent despite such adversarial manipulations. In clone detection, a robust model should consistently identify clones even after minor, semantic-preserving transformations, such as variable renaming or control structure rewriting. In just-in-time defect prediction, robustness implies that modifications—such as adding dead code or removing redundant lines—should not cause the model’s prediction to flip from clean to buggy, or vice versa. Formally, for a model $f:X\rightarrow Y$, adversarial robustness implies that for any small perturbation $\delta$ bounded by a constraint (e.g., $|\delta|\leq\epsilon$ for some small $\epsilon>0$), the model satisfies $f(x)=f(x+\delta)$ for most inputs $x\in X$ [18, 63, 64].
@@ 2 Background
Machine Learning Explainability: Machine Learning (ML) explainability refers to understanding, interpreting, and articulating a machine learning model’s internal mechanisms, reasoning processes, and outputs. In other words, explainability seeks to provide human-interpretable insights into why a model arrives at a particular prediction or decision [58]. Common approaches to ML explainability include feature importance analysis, surrogate modeling (e.g., LIME, SHAP), attention mechanisms, and rule extraction. These methods help practitioners and researchers interpret model outputs and identify biases, errors, and potential vulnerabilities in the models, including their susceptibility to adversarial attacks.
3 Research Methodology
@@
Figure 1: Our proposed workflow for evaluating the robustness of ML models under adversarial attacks. Step 1 includes the dataset selection and preprocessing methods. Step 2 involves training ML models, while Step 3 involves selecting ML explainability techniques. Step 4 then describes how we perform adversarial attacks using ML explainability, and Step 5 investigates the robustness of ML models under adversarial attacks.
3.1 Dataset Selection and Preprocessing
@@
This study aims to evaluate the robustness of ML models by generating adversarial examples through feature-space manipulation. We meticulously select and process the datasets as follows:
3.1.1 Dataset Selection
@@
This study investigates the robustness of ML models against adversarial attacks across various software analytics tasks, including defect prediction, clone detection, and code review comment classification. Additionally, our study shifts the focus of attack from the input-space to the feature-space when generating adversarial examples. Thus, we aim to select software analytics tasks where features extracted from the input (e.g., source code) are utilized to perform the tasks of interest. For example, in cross-language clone detection, various features such as the ‘Number of Variables Declared,’ ‘Total Number of Operators,’ ‘Number of Arguments,’ ‘Number of Expressions,’ ‘Total Number of Operands,’ ‘Number of Loops (for, while),’ ‘Number of Exceptions Thrown,’ ‘Number of Exceptions Referenced,’ and ‘McCabe’s Cyclomatic Complexity’ were extracted first, and ML models were then trained based on these extracted features [1]. Consequently, we select six different datasets: three for defect prediction, two for clone detection, and one for code review comment classification, as the targets for adversarial attacks, where features are extracted from the input (e.g., source code) and stored in tabular format.
Table 1: Summary of datasets used in this study
Task Dataset # Data Points # Features
JIT-defect Prediction Cross Project 30,000 14
Java Project 25,000 27
Postgres 20,431 14
Clone Detection CLCDSA 30,000 18
BigCloneBench 5,46,000 6
Code Review Comments Classification Code Review 1,100 15
@@
Table 1 summarizes the key characteristics of the datasets used in this study. For JIT defect prediction, we utilize a dataset for cross-project mobile apps [13] and ten defect prediction datasets from various Java projects [14]. Additionally, we select the Postgres dataset [15] from six subject systems for the Desktop application. The code review comment dataset [39] comprises 15 features and approximately 1,100 data points for detecting useful comments. Finally, we extend our experimental analysis by incorporating the cross-language clone dataset used by Nafi et al. [1] and the BigCloneBench (BCB) dataset [50]. The selected cross-project mobile apps dataset contains 14 features and approximately 30,000 data points, while different Java projects have datasets with 65 features and approximately 25,000 data points. The CLCDSA dataset consists of 18 features and approximately 30,000 data points. Feng et al. [2] extracted six features from the BCB dataset based on some similarity metrics. In our study, we also use these six features to train the ML models. The preprocessed dataset ended up with six features and approximately 546K samples. Please refer to these articles [50, 15, 13, 1, 39] for details about the features of each dataset.
3.1.2 Feature Engineering